a) Define elasticity of a matter.



b) Define stress and strain. Explain different types of stress and strain. 1+4



c) The radius of a silver bar is 30 ๐‘๐‘š. What force is required to stretch the bar by 12.7% of its length where Youngโ€™s modulus is ๐‘Œ = 7.25 ร— 1010 ๐‘๐‘šโˆ’2 .

1
Expert's answer
2021-12-07T09:52:24-0500

Elasticityย is the property of solid materials to return to their original shape and size after the forces deforming them have been removed.ย 

What is Stress?

The force applied per unit area in mechanics is known as stress.

What is Strain?

Strain is the ratio of the amount of deformation experienced by the body in the direction of force applied to the initial sizes of the body.ย 

Types of Stress

Normal Stress: The restoring force per unit area perpendicular to the body surface is known as the normal stress. It differentiated into two types: tensile and compressive stress.ย 

ย Tangential Stress: It is called tangential stress when the elastic restoring force acts parallel to the surface area.

Types of Strain

Longitudinal Strain: The strain produced on the body due to the deforming force, which leads to change in only the length of the object is known as longitudinal or the tensile strain.ย 

Volumetric Strain: This is the strain produced on the body due to the deforming force, which leads to only the change in volume of the object.

Shear Strain: Due to the tangential stress, an angle tilt is caused in the body; this is known as shear strain.


"7.25\cdot10^{10}=\frac{F}{\pi(0.3)^20.127}\\F=2.60\cdot10^{9}\ N"